OpenStudy (zaxoanl):

can someone show me how to find the length of vector v whose initial point is (4,0) and terminal point is (-2,1)

OpenStudy (3mar):

Could I help you?

OpenStudy (3mar):

Are you familiar with that? @zaxoanl

OpenStudy (zaxoanl):

thank you @3mar

OpenStudy (3mar):

Don't mention it! You are welcome! What matters is that you got the idea! Did you?

OpenStudy (zaxoanl):

i have another question Find the equation of the line that passes through (-4,5,2) and is parallel to both the xz-plane and yz-plane. i got y-5=0 y=5 x+4=0 x=-4

OpenStudy (zaxoanl):

the equation would be (x+4)+(y-5)+(z-2)=0 right?

OpenStudy (3mar):

It would better to start your second question in another new post if you don't mind!